OpenStudy (anonymous):

a ranger catches 319 deer in a game preserve tags them then releases them. Later 84 deer are caught and it is found that 32 of them are tagged. Estimate how many deer are in the game preserve (round to the nearest integer.)

OpenStudy (anonymous):

837 deer in the preserve :)

OpenStudy (anonymous):

thank you so much how would I have firgured that answer?

OpenStudy (anonymous):

nevermind I figured out how you did it thanks again so so much

OpenStudy (anonymous):

32/84*100=38.1% 319=38.1% 100%=? 1%=319/38.1%=8.37 8.37*100=837 deer :)